OpenStudy (anonymous):

Please Help! -4.5x - 2y = -12.5 3.25x - y = -0.75 I don't even know where to begin

OpenStudy (secret-ninja):

http://www.wolframalpha.com This website allows you to input any question for help.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

if you take the second equation and isolate y, you get: y=3.25x+0.75 now you can place it in the first equation and solve for x: -4.5x -2(3.25x+0.75)=-12.5

OpenStudy (anonymous):

-4.5x-6.5x-1.5=-12.5 -11x=-12.5+1.5 -11x=-11 /: -11 x=1